OpenStudy (anonymous):

It says to graph the following equations but i don't know how to solve the equations to get the coordinates in order to graph: Y=-1/3x

OpenStudy (anonymous):

okay so the point-slope form of any equation/graph is y = mx + b you don't have to solve anything, silly :) m = slope = rise/run b = y-intercept (where the line crosses the y-axis) so, this equation says: y = (-1/3)x + 0 right? so... m= -1/3 rise= -1 (down one) run = +3 (right three) y-intercept = 0 So....start with your y-intercept.

OpenStudy (anonymous):